how many NA+ and Cl
- ions are present in one unit cell of nacl???
 
 

Grade:12

2 Answers

Krishna Murmu
7 years ago
4 Na+ and 4 Cl- ions are present in unit cell. NaCl has rock salt structure. Where Cl- occupies FCC cubic lattice sites and Na+ occupies octahedral voids.
Me
7 years ago
NaCl is a kind of FCC unit cell
 
*Anions are involved in closepacking. So Cl^-ions will be present at all the corners as well as at centre of each face
       Contribution of one anion at corner= 1/8
         Total contribution = 1/8 × no. of corners = 1/8 × 8= 1
       Contribution of one anion at face centre = ½
          Total contribution = 1/2×no. Of faces = 1/2×6= 3
So, total number of anions =1+3 = 4
 
* Cations move to the voids . So Na+ ions will be present in the octahedral voids ( they don't move to tetrahedral voids as they are larger in size than the tetrahedral voids).
       No. Of octahedral voids = no. Of constituent particles = 4
So, we have total 4 Na + ions in a unit cell
 
=>  4 Na+ ions and 4 Cl- ions.